# Homophily-Driven Sanitation View for Robust Graph Contrastive Learning

### v2

@@ IV-A Threat Model
We consider an adversarial environment involving two opposing roles. A graph attacker seeks to inject topology noise into the clean graph to intentionally reduce the quality of the node embeddings produced by GCL, while the defender aims to improve embeddings to enhance the performance of downstream tasks. In practical scenarios, the defender typically collects data from the environment and constructs a graph, after which the GCL models transform the graph data into low-dimensional node embeddings for ease of analysis. However, the graph attacker may interfere with the data processing procedure and create a poisoned graph to degrade the performance of graph-based tasks.
Faced with attacks against GCL, our primary goal is to design a robust GCL model for the defender. We assume that the defender does not know the detailed generation mechanism of the attack graph as well as the strength of the attack. The defender only has access to a poisoned graph and importantly, any label information is not available to the defender.
IV-B Problem of Defense
In this paper, our primary goal is to develop an adversarial robust GCL model that can generate high-quality embeddings from poisoned graphs. While the model robustness is often measured by its performance on downstream tasks, we go one step deeper to analyze the robustness through an information perspective. Specifically, we adopt a metric from the literature to measure robustness:
Definition 1 (GRV [29, 30]).
Definition 2 (Defense of GCL).
V Vulnerability Analysis of GCL
In this section, we provide an in-depth analysis, both empirically and theoretically, of how existing attacks would affect GCL through the lens of an important property of graphs. These explorations of the vulnerabilities of GCL provide strong insights for our design of the robust GCL.
V-A Empirical Analysis of Attack Effects
A fundamental property that guides GCL is homophily, which states that the nodes connected by edges in the graph tend to be more similar. In fact, the majority of the graphs observed in the real world are homophilous [8]. Previous studies have designed various metrics to measure the degree of homophily of a graph. A primary metric is to calculate the percentage of the intra-class links in a graph:
Definition 3 (Graph homophily on label space [8]).
The graph homophily based on node labels is given by:
where $y_{{i}}$ is the label of node $v_{i}$, $\mathcal{E}$ is the link set.
Besides using node labels, the graph homophily can also be measured by the Euclidean distance between the attribute vectors of the connected nodes [31, 16]:
Definition 4 (Graph homophily on feature space [16]).
The graph homophily based on node attributes is given by:
A higher $h_{y}$ (or a lower $\delta_{x}$) indicates that the graph tends to be more homophilous.
Previous studies [32] have revealed that attacks against GNNs (e.g., Mettack [12]) tend to inject inter-class edges into the clean graph to downgrade the performance of GNNs. We then conduct comprehensive experiments to verify whether attacks against GCL models have similar properties. Specifically, we test an existing attack method (termed CLGA [13]) against a particular GCL model (GCA [4]) on the Cora dataset as an example. We also test Mettack on GCL model. We firstly train CLGA and Mettack (Mettack can access to the data split) on the clean graph and generate the corresponding poisoned graph with multiple attacking powers ranged in $\{5\%,10\%,15\%,20\%\}$. Then, we feed the GCL model with the poisoned graph to evaluate the performance on node classification. Fig. 1 presents the relationship between the semi-supervised node classification accuracy of GRACE [3] (a typical GCL baseline) and the graph homophily metrics ($h_{y}$ and $\delta_{x}$) on poisoned graphs with varying attacking powers. It is clearly observed that the accuracy is positively correlated with $h_{y}$ and is negatively correlated with $\delta_{x}$. That is, both attacks will significantly decrease the homophily of the graph.
(a) Mettack
(b) CLGA
Table I: Number of inserted intra-class links, inserted inter-class links, deleted intra-class links and deleted inter-class links for Mettack and CLGA on Cora dataset.
Attack Attack power $5\%$ $10\%$ $15\%$ $20\%$
Mettack #Inserted intra-class links $3$ $9$ $14$ $25$
#Inserted inter-class links $248$ $486$ $705$ $914$
#Deleted intra-class links $2$ $11$ $38$ $68$
#Deleted inter-class links 00 00 $3$ $6$
CLGA #Inserted intra-class links $18$ $35$ $59$ $77$
#Inserted inter-class links $232$ $465$ $690$ $912$
#Deleted intra-class links $3$ $5$ $10$ $21$
#Deleted inter-class links 00 $1$ $1$ $3$
Figure 2: A GCL framework with homophily-driven learnable sanitation view.
In addition, Tab. I shows the number of inserted intra-class links, inserted inter-class links, deleted intra-class links, and deleted inter-class links for Mettack and CLGA at different attacking powers. The results indicate that, for both CLGA and Mettack, the majority of manipulations involve inserting inter-class links into the clean graph. This finding is consistent with the phenomenon presented in Fig. 1. All these experiment results demonstrate that attacks against GCL take effect by reducing the homophily levels of graphs.
V-B Theoretical Analysis
In this section, we provide the theoretical analysis of how the attacks can negatively impact the learning of the GCL model using mutual information as the bridge. For ease of proof, we first introduce a lemma:
@@ V-B Theoretical Analysis
The two attacks, Mettack and CLGA, maliciously degenerate the GCL’s performance by minimizing the mutual information between the graph and its representations.
@@
Based on Eqn. 11 and 12, we have
$\begin{split}I(G;\mathbf{Z})=&H(G)-\mathcal{L}_{CE}+D_{KL}(p_{YZ}\|p_{Y}p_{Z})+D_{KL}(p_{Y}\|p_{Z})\\ +&H(\mathbf{Y}|G,\mathbf{Z})-H(G|\mathbf{Y},\mathbf{Z}).\end{split}$ (13)
Based on Eqn. 13, we observe that the mutual information $I(G;\mathbf{Z})$ is negatively correlated with $\mathcal{L}_{CE}$. Since Mettack can increase the $\mathcal{L}_{CE}$:
@@
then, Eqn. 9 also holds for the poisoned graph generated by Mettack. ∎
@@
Theorem 1 demonstrates that by maliciously altering the graph topology, attacks can decrease the mutual information between the graph data and its representations. On the other hand, GCL’s high-quality node representations are derived from maximizing this mutual information. Therefore, the GCL model may only be able to optimize the mutual information to a sub-optimal level, resulting in the production of low-quality embeddings.
VI Defense Methodology
In this section, we first provide an overview of our framework to clearly describe the whole picture and then clarify our homophily-driven learnable sanitation view for robust graph contrastive learning.
VI-A Overview
@@
In this paper, we propose the homophily-driven learnable sanitation view for enhancing the robustness of unsupervised graph contrastive learning. The proposed model’s overall framework is illustrated in Fig. 2. Similar to the vanilla GCL framework, our model comprises a data augmentation phase and a contrastive learning phase. In the first phase, we introduce a homophily-driven learnable sanitizer that stochastically generates a sanitation view. This learnable sanitizer is parameterized by the link removal probability matrix. In the second phase, the sanitizer generates a sanitation view and contrasts it with the second augmented view, which is produced by random link removal. We then feed these two views to a shared GNN encoder and generate two node features for contrast. Finally, we optimize a homophily-driven objective in an end-to-end manner based on the two node features, updating the link removal probability matrix and the GNN encoder’s parameters via gradient descent.
VI-B End-to-End Design of Learnable Sanitizer
In this section, we introduce our design and implementation of the learnable sanitizer, which can generate sanitation views for GCL. In particular, this sanitizer is seamlessly embedded into the GCL framework and is jointly trained with the contrastive learning process.
VI-B1 Edge-dropping sanitizer
@@
Our design of the sanitizer relies on a stochastic edge-dropping scheme that aims to sanitize the inter-class links injected by the attacker. This design is based on the preliminary experiments (Section V-A) that existing attacks tend to insert inter-class links into the clean graph. Specifically, we use an edge vector $\mathbf{E}$ to represent all the existing edges in a graph. Then, we define a manipulation vector $\mathbf{M}=\{0,1\}^{|\mathbf{E}|}$ that follows the Bernoulli distribution parameterized by $\mathbf{P}$; that is $\mathbf{M}\sim Ber(\mathbf{P})$. For a particular entry $\mathbf{M}_{e}$, we will remove the corresponding edge $e$ if $\mathbf{M}_{e}=1$, which is controlled by the probability $\mathbf{P}_{e}$. Now, the sanitizer can be regarded as a mapping $\mathcal{T}_{\mathbf{P}}(\mathbf{E})\rightarrow\mathbf{E}_{\mathcal{S}}$ parameterized by $\mathbf{P}$, which takes a graph with edge vector $\mathbf{E}$ as input and produce a sanitized view $\mathbf{E}_{\mathcal{S}}$. Formally, we have:
@@
where ($\odot$($\cdot$)⊙ ( ⋅ ) is the element-wise product operation, $\mathbf{1}_{|\mathbf{E}|}$ is an all-one vector with length equal to the edge number ${|\mathbf{E}|}$, $\mathbf{P}=\{p_{e}\}_{e=1}^{|\mathbf{E}|}$ is the parameters of the sanitizer.
@@
VI-B2 End-to-end learning
The key challenges to learning $\mathcal{T}_{\mathbf{P}}$ are two-fold: choosing the proper objective to guide learning, and integrating learning into the framework of GCL. In this paper, we design the objective function as follows (while deferring the explanation to the next section):
@@ VI-B2 End-to-end learning
where the first view $\mathbf{E}_{\mathcal{S}}$ is the sanitation view generated from the sanitizer, the second view $t_{\mathcal{R}}(\mathbf{E})$ is generated from typical random augmentation scheme $\mathcal{T}_{R}$ that will randomly drop edges with a fixed probability, $\operatorname{Tr}(\cdot)$ represents the trace of a matrix and $\mathbf{S}_{1}$ is a constrained space of the sanitation operation with $\epsilon$ controlling the sanitation degree:
@@
In our design, the training objective of the sanitizer is embedded into that of the GCL. That is, we train the sanitizer together with the contrastive learning, using a hyperparameter $\eta$ to control the relative importance of optimizing the infoNCE loss and the graph homophily $\delta_{x}$. Then, the learnable sanitizer sample the sanitation view $\mathbf{E}_{\mathcal{S}}$ based on Eqn. 15 for each iteration. After that, the shared GNN encoder $f_{\theta}(\cdot)$ converts the graph data from the two views to low-dimensional node embedding matrices and obtain the training loss $\mathcal{L}$. During the backward pass, we compute the gradient of $\mathcal{L}$ with respect to the parameters $\mathbf{P}$ and $\theta$ and update them via gradient descent.
VI-B3 Insights behind our design
The objective function of our robust model in Eqn. 16 contains two vital components: the infoNCE loss over two views and the graph homophily measurement $\delta_{x}=\operatorname{Tr}(\mathbf{X}^{\top}\mathbf{LX})$. In Sec. V-B, it was (Theorem 1) shown that the attacks would degrade the consistency (mutual information) between the graph and its representations, leading to a decrease in the quality of node embeddings. Therefore, an effective robust model should restore the properties distorted by malicious topology manipulations, acting as a countermeasure against the graph attacker. To achieve this, the sanitizer should be designed in a way that the mutual information between graph and representation is preserved. As the infoNCE objective is a common approximation to the intractable mutual information, we optimize the sanitation view to minimize the infoNCE loss (i.e., maximize the mutual information). There are other graph sanitation frameworks for GNNs, such as those proposed in [34, 35], which adopt the cross-entropy loss as the objective for sanitation. However, these are infeasible in an unsupervised learning setting as they require access to node labels for training. Nevertheless, we want to emphasize that we are the first to theoretically describe the inherent relationship between the objective of the GCL model and the graph attacks, which provides strong insights for learning the sanitation view.
Furthermore, as discussed in Sec. V-A, it has been observed that attacks (Mettack and CLGA) tend to add heterophilous links, which can decrease the homophily degree of the graph (e.g., increasing $\delta_{x}$). Based on this observation, we choose the graph homophily metric $\delta_{x}$ as a second supervision term to sanitize the poisoned graph to obtain the sanitation view. Compared to $h_{y}$, which requires node labels, $\delta_{x}$ does not, and is thus suitable for unsupervised learning. In Sec. VII, we provide a comprehensive ablation study to verify the significance of these two components.
VI-C Implementation
Now, we introduce several techniques to efficiently solve the optimization problem (16).
VI-C1 Gumbel-Softmax Re-parametrization
To optimize the parameters $\mathbf{P}$ via gradient descent, it is necessary to tackle the in-differentiation of the sampling procedure in Eqn. 15 during the backward pass. To tackle this problem, we utilize the Gumbel-Softmax re-parametrization technique [19] to relax the discrete Bernoulli sampling procedure to a differentiable Bernoulli approximation. Using the relaxed sample, we implement the straight-through estimator [20] to discretize the relaxed samples in the forward pass and set the gradient of discretization operation to $1$, i.e., the gradients are directly passed to the relaxed samples rather than the discrete values. Hence, we reconstruct the mapping from $\mathbf{P}$ to $\mathbf{M}$ as:

### v3

@@ IV-A Threat Model
We consider a system consisting of two parties: an attacker and a defender. Specifically, the attacker can manipulate the graph data to degenerate the graph-based learning models deployed by the defender. Conversely, the defender’s objective is to recover the quality of the graph representation learning given the poisoned graph to improve node classification performance. To be concrete, we summarize the attacker’s goal (Mettack and CLGA as exemplar), knowledge, and capability as follows:
  • Attacker’s goal: The attacker’s goal is assumed to be decreasing the classification accuracy of a node classification problem achieved after training on poisoned graph data manipulated by the attacker.
  • Attacker’s knowledge: The attacker can have different levels of knowledge of the data, model, and model parameters. For Mettack and CLGA, the attacker has no knowledge about the model with its parameters. In the meanwhile, the attacker can observe all nodes’ attributes, the graph structure. The main difference between Mettack and CLGA is that Mettack can observe the training labels. However, CLGA cannot query the label information of the graph data since it is an unsupervised attack method.
  • Attacker’s capability: In order to achieve unnoticeable attacks, Mettack and CLGA both impose a budget constraint $\Delta$ to restrict the number of perturbations on the topology space of the graph data, i.e., $\frac{1}{2}\|\mathbf{A}-\mathbf{A}^{p}\|\leq\Delta$, where $\mathbf{A}^{p}$ is the poisoned adjacency matrix.
  • Defender’s knowledge: In reality, the defender cannot acquire the attacking scenarios (such as the attack type, attack degree, and victim nodes or edges, etc.), and the label information of the graph data. However, the defender can have full knowledge of the given graph data’s information (the node attribute matrix and adjacency matrix) as well as the model’s information (model structure and model parameters).
IV-B Disturbance on Mutual Information
Figure 1: The blue and green circles represent the information entropy of the graph $G$ and embedding $\mathbf{Z}$, with the shaded area indicating the mutual information.
The powerful graph representation learning of the GCL framework highly relies on capturing the mutual information between the graph and its representations in low-dimensional space. To explore the failure of the GCL framework under the adversarial attack scenario, we start by analyzing the vulnerability of the GCL framework from an information theoretical perspective. For graph contrastive learning, a common way to curve the intractable mutual information is to utilize the InfoNCE object:
@@ IV-B Disturbance on Mutual Information
The two attacks, Mettack and CLGA, maliciously degenerate the GCL’s performance by diminishing the mutual information between the graph and its representations.
@@
Based on Eqn. 6 and 7, we have
$\begin{split}I(G;\mathbf{Z})&=H(G)-\mathcal{L}_{CE}+D_{KL}(p_{YZ}\|p_{Y}p_{Z})\\ &+D_{KL}(p_{Y}\|p_{Z})+H(\mathbf{Y}|G,\mathbf{Z})-H(G|\mathbf{Y},\mathbf{Z}).\end{split}$ (8)
Based on Eqn. 8, we observe that the mutual information $I(G;\mathbf{Z})$ is negatively correlated with $\mathcal{L}_{CE}$. Since Mettack can increase the $\mathcal{L}_{CE}$:
@@
then, Eqn. 4 also holds for the poisoned graph generated by Mettack. ∎
((a))
((b))
Figure 2: $I_{InfoNCE}$ for Cora dataset attacked by Mettack and CLGA with varying attacking powers.
@@
Theorem 2 demonstrates that by maliciously altering the graph topology, attacks can degrade the mutual information estimation between the graph data and its representations. On the other hand, GCL’s high-quality node representations are derived from maximizing this mutual information. Therefore, the GCL model may only be able to optimize the mutual information to a sub-optimal level, resulting in the production of low-quality embeddings. We further conduct empirical studies to verify whether the structural attacks may maliciously influence mutual information. We first implement Mettack and CLGA individually to poison the clean graph (Cora [15]) under different attacking scenarios. Next, we train a simple GCL model, i.e., GRACE [3] on the poisoned graphs and report the converged InfoNCE object in Fig. 2. The results demonstrate that the structural attacks on GRACE indeed degrade the mutual information estimation between the graph and its representations.
IV-C Disturbance on Graph Homophily
It has been widely explored that structural attacks against GNNs (Mettack) will diminish the homophily level of the clean graphs [33, 32, 38, 12, 34]. In this paper, we also explore whether the above-mentioned phenomenon is a common principle for graph structural attacks, including Mettack and CLGA. In fact, the graph homophily is defined as:
Definition 3 (Graph homophily on label space [10]).
The graph homophily based on node labels is given by:
where $y_{{i}}$ is the label of node $v_{i}$, $\mathcal{E}$ is the link set.
Besides using node labels, the graph homophily can also be measured by the Euclidean distance between the attribute vectors of the connected nodes [37, 32]. It is common to use a trace formula to reformulate such distances in matrix format. We then present another metric on graph homophily as follows:
Definition 4 (Graph homophily on feature space [32]).
The graph homophily based on node attributes is given by:
A higher $h_{y}$ (or a lower $\delta_{x}$) indicates that the graph tends to be more homophilous. We conduct empirical studies to verify whether structural attacks will diminish the graph homophily is a common principle for Mettack and CLGA. Fig. 3 presents the relationship between the node classification accuracy of GRACE [3] and the graph homophily metrics ($h_{y}$ and $\delta_{x}$) on poisoned graphs with varying attacking powers. It is clearly observed that the accuracy is positively correlated with $h_{y}$ and is negatively correlated with $\delta_{x}$. That is, both attacks will significantly diminish the graph homophily.
((a))
((b))
Overall, the findings in Sec. IV-B and IV-C that adversarial attacks will provide malicious effects to the mutual information and the graph homophily can serve as strong insights for us to design the defense strategy for the GCL framework. That is, restoring the diminished mutual information and graph homophily is promising to sanitize the malicious effects and thus achieve adversarial robustness.
V Defense Methodology
Figure 4: The overall architecture of GCIR.
V-A Overview
@@
The proposed model’s overall framework is illustrated in Fig. 4. The model comprises a data augmentation phase and a contrastive learning phase. In the first phase, we introduce a learnable sanitation view to restore the diminished mutual information and graph homophily. In the second phase, the sanitation view is contrasted with the second augmented view, which is produced by random link removal and feature masking. Finally, we optimize a cross-view contrastive learning objective and the learnable sanitation view in an end-to-end manner. In this way, the learned node embeddings will restore the diminished mutual information and achieve adversarial robustness.
V-B Sanitation View Generation
@@
As previously mentioned in Sec. I, we endeavor to achieve the robustness of GCL via restoring the contaminated mutual information after attacks. Unlike the traditional adversarial training method that achieves robustness by further “polluting” the poisoned graph data, we instead choose to “sanitize” the poisoned graph data to rectify the mutual information distortion caused by the attacks during the data augmentation phase. To achieve this goal, we introduce the sanitation view with a stochastic edge-dropping scheme to sanitize malicious effects provided by attacks. To be concrete, the stochastic edge-dropping scheme of the sanitation view can adaptively restore the mutual information during graph data augmentation and thus achieve robustness. Specifically, we use an edge vector $\mathbf{E}$ to represent all the existing edges in a graph. Then, we define a manipulation vector $\mathbf{M}=\{0,1\}^{|\mathbf{E}|}$ that follows the Bernoulli distribution parameterized by $\mathbf{P}$; that is $\mathbf{M}\sim Ber(\mathbf{P})$. For a particular entry $\mathbf{M}_{e}$, we will remove the corresponding edge $e$ if $\mathbf{M}_{e}=1$, which is controlled by the probability $\mathbf{P}_{e}$. Now, the sanitation view can be regarded as a mapping $\mathcal{T}_{\mathbf{P}}(\mathbf{E})\rightarrow\mathbf{E}_{\mathcal{S}}$ parameterized by $\mathbf{P}$, which takes a graph with edge vector $\mathbf{E}$ as input and produces a sanitized view $\mathbf{E}_{\mathcal{S}}$. Formally, we have:
@@
where $\odot(\cdot)$ is the element-wise product operation, $\mathbf{1}_{|\mathbf{E}|}$ is an all-one vector with length equal to the edge number $|\mathbf{E}|$}, $\mathbf{P}=\{p_{e}\}_{e=1}^{|\mathbf{E}|}$ is the parameters of the sanitizer.
@@
V-B1 Learning of Robust GCL
The key challenges to learning $\mathcal{T}_{\mathbf{P}}$ are two-fold: choosing the proper objective to guide learning, and integrating learning into the framework of GCL. We design an objective function that contains two components, i.e., InfoNCE loss and feature smoothness. Intuitively, forcing the learning of the sanitation view to decrease the InfoNCE loss and feature smoothness can restore the diminished mutual information and graph homophily during training. Specifically, the objective function is :
@@ V-B1 Learning of Robust GCL
where the first view $\mathbf{E}_{\mathcal{S}}$ is the learnable sanitation view, the second augmented view $t_{\mathcal{R}}(\mathbf{E})$ is generated from typical random augmentation scheme $\mathcal{T}_{R}$ that will randomly drop edges and feature masking with a fixed probability, $\operatorname{Tr}(\cdot)$ represents the trace of a matrix and $\mathbf{S}_{1}$ is a constrained space of the sanitation operation with $\epsilon$ controlling the sanitation degree:
@@
In our design, the training objective of the sanitation view is embedded into that of the GCL. That is, we train the learnable sanitation view together with the contrastive learning, using a hyperparameter $\eta$ to control the relative importance of optimizing InfoNCE loss and the graph homophily $\delta_{x}$. Then, the learnable sanitation view samples $\mathbf{E}_{\mathcal{S}}$ based on Eqn. 12 for each iteration. After that, the shared GNN encoder $f_{\theta}(\cdot)$ converts the graph data from the two views to low-dimensional node embedding matrices and obtains the training loss $\mathcal{L}$. During the backward pass, we compute the gradient of $\mathcal{L}$ with respect to the parameters $\mathbf{P}$ and $\theta$ and update them via gradient descent. Consequently, the sanitation view can restore the diminished mutual information guided by the sanitation view’s objection.
V-B2 Learning of Stochastic Edge-dropping Scheme
However, the above-mentioned learning scheme still encounters a unique challenge: It is difficult to directly optimize the stochastic edge-dropping scheme of the sanitation view due to the undifferentiable sampling procedure. To tackle this issue, we refer to the Gumbel-Softmax re-parametrization technique [50] to relax the discrete Bernoulli sampling procedure to a differentiable Bernoulli approximation. Using the relaxed sample, we implement the straight-through estimator [51] to discretize the relaxed samples in the forward pass and set the gradient of discretization operation to $1$, i.e., the gradients are directly passed to the relaxed samples rather than the discrete values. Hence, we reconstruct the mapping from $\mathbf{P}$ to $\mathbf{M}$ as:
@@ V-B2 Learning of Stochastic Edge-dropping Scheme
where $\lfloor\cdot\rceil$ is the rounding function, $g\sim Gumbel(0,1)$ is a standard Gumbel random variable with zero mean and the scale equal to $1$. Then, we can reformulate Eqn. 12 to:
Formally, to optimize the parameters $\mathbf{P}$ under the constraint $\mathcal{S}_{1}$ defined in Eqn. 16, we deploy the projection gradient descent to optimize the relaxed constrained optimization problem introduced in Eqn. 16:
at the $t$-th iteration, where $\alpha$ is the learning rate of the projection gradient descent, $\nabla_{\mathbf{P}}\mathcal{L}(t_{1}(\mathcal{G}),t_{2}(\mathcal{G}),\theta)$ denotes the gradient of the loss defined in Eqn. 13, $\prod_{\mathcal{S}_{1}}(\cdot)$ is the projection operator to project the updated parameters to satisfy the constraint. Referring to [21], the projection operator $\prod_{\mathcal{S}_{1}}(\cdot)$ has the closed-form solutions:
where $\prod_{[0,1]}[\mathbf{P}]$ clips the parameters vector $\mathbf{P}$ into the range $[0,1]$. The bisection method [52] is used over $\mu\in\{\min(\mathbf{P}-\mathbf{1}_{E}),\max(\mathbf{P})\}$ to find the solution to $\sum_{e}\prod_{[0,1]}[\mathbf{P}-\mu\mathbf{1}_{E}]=\epsilon$ with the convergence rate $\log_{2}[(\max(\mathbf{P})-\min(\mathbf{P}-\mathbf{1}_{E}))/\xi]$ with $\xi$-error tolerance [53]. We present the details of algorithm in Alg. LABEL:alg-RGCL.
Algorithm 1 GCIR
V-C Unsupervised Tuning Strategy
Strictly speaking, the whole training procedure of the GCL models is label-free, that is, the data analyst cannot query the label’s information from the downstream tasks to supervise the training of GCL, including the parameter tuning phase. Hence, it is vital to craft a totally unsupervised tuning strategy to select the best hyperparameters. We herein concentrate on designing a new metric in an unsupervised manner without additional training or fine-tuning.
In particular, normalized cut loss [42] is an unsupervised loss function that counts the fraction of the inter-class links for a given graph partition result. Thus, a good graph partition should obtain a lower normalized cut loss, i.e., the majority of the links connect two nodes with the same class. The normalized cut loss is defined as: